Four charges, q1 =5.00 x10-7C,   q2 = -3.00 x10-7 C,   q3 = -2.00 x10-7C,   q4 = 6.00 x10-7C, are situated at the corners of a square of side 4.00m. Find the potential at the center of the square?



1
Expert's answer
2022-02-19T04:55:58-0500

Given:

"q_1=5.00*10^{-7}\:{\rm C},\quad q_2=-3.00*10^{-7}\:\rm C"

"q_3=-2.00*10^{-7}\:{\rm C},\quad q_4=6.00*10^{-7}\:\rm C"

"r=\frac{\sqrt{2}}{2}*4.00\:\rm m=2.83\: m"


The net potential at the center of square

"V=V_1+V_2+V_3+V_4"

"V=\frac{k}{r}(q_1+q_2+q_3+q_4)"

"V=\frac{9*10^9}{2.83}(5.00-3.00-2.00+6.00)*10^{-7}=1910\:\rm V"